### v2

# Style2Code: A Style-Controllable Code Generation Framework with Dual-Modal Contrastive Representation Learning Style2Code: A Style-Controllable Code Generation Framework with Dual-Modal Contrastive Representation Learning

Abstract
Controllable code generation aims to produce source code that adheres to a specified stylistic pattern while maintaining its functional correctness. However, existing approaches often rely on implicit style cues or handcrafted templates, which limit their flexibility and generalizability. In this paper, we propose Style2Code, a novel multimodal framework for style-controllable code generation. Our approach explicitly encodes code style into a continuous vector representation that captures fine-grained stylistic attributes, such as naming conventions, indentation patterns, and structural layout. This style vector is then concatenated with source code tokens and fed into a pre-trained language model (e.g., Flan-T5), enabling the decoder to generate code that reflects the target style. Experimental results on multiple benchmark datasets demonstrate that Style2Code significantly improves both stylistic alignment and generation quality, outperforming state-of-the-art baselines on BLEU, ROUGE, and Code Style Similarity (CSS) metrics. To the best of our knowledge, Style2Code is the first model to achieve explicit and user-controllable code style transfer through continuous vector conditioning in a dual-modal generation setting.
The source code and data sets are publicly available to facilitate further research:
Style2Code: A Style-Controllable Code Generation Framework with
Dual-Modal Contrastive Representation Learning
1 Introduction
1.1 Problem
In the daily use of AI models for code generation, it is often observed that the generated code may not fully meet the user’s requirements in terms of style due to the influence of training data and user prompts. Additionally, the code style produced by different models may also vary slightly. Without providing examples of similar types of code, specifying a particular style in the prompt, or providing detailed algorithm requirements, the model’s output often contains an element of randomness. We can categorize the style variation issues in code-generative AI into several types: (1) differences in style output by the same model for identical prompts at different times; (2) style variations between different models at the same time for identical prompts; (3) differences in output from the same model at the same time for prompts with similar structure but different style requirements. The code style itself also encompasses multiple aspects: (1) naming differences, such as variations in variable names; (2) functional differences, where the same code may implement functionality in varied ways; (3) structural or functional discrepancies arising from a lack of context, leading to incomplete structure or function implementation.
The same prompt can lead to different code styles when generated by different models; In simple terms,Even for the same model, repeated generations may show slight stylistic variations (despite similar structure).
1.2 Background
Personalized code generation has become an important area in automated code synthesis. MPCoder advances this field by incorporating both explicit and implicit style representations for multi-user scenarios. However, it faces key limitations: it cannot effectively generalize implicit stylistic cues embedded in code semantics and relies on fixed style encodings and handcrafted metrics, limiting adaptability.
Additionally, MPCoder demands large user-specific datasets, creating scalability issues—each new user requires retraining, leading to cold-start problems and high maintenance costs. This hinders practical deployment in dynamic environments with limited training data.
While works like Choi et al. (2023) explore stylistic clustering for authorship attribution using LLMs, they focus on identification, not controllable generation. Thus, there remains a clear need for a scalable, generalizable framework that enables fine-grained, user-controllable code style generation with minimal data and no retraining.
1.3 Contributions
Our main contributions are summarized as follows:
(1) We design a fine-grained explicit style encoder that directly extracts interpretable stylistic features—such as indentation schemes, whitespace alignment, loop nesting depth, and structural repetition patterns—from raw source code, without relying on manual annotations or handcrafted templates.
(2) We introduce a contrastive alignment framework that semantically aligns code embeddings with their corresponding style vectors, enabling the model to generalize to unseen styles while preserving controllability during generation.
(3) We propose a two-stage training paradigm that first trains the style encoder independently to obtain stable and accurate style representations, and then injects these representations into the code generator for style-controlled generation. This decoupled training strategy improves overall training stability and significantly enhances the effectiveness of style transfer.
(4) Furthermore, we formulate the style transfer task as a multimodal learning problem, and explore its potential in the context of personalized code generation, enabling flexible and user-controllable stylistic variation.
2 Method
We define the personalized code generation task as amultimodal generation problem, where the goal is to generate code that both retains the original functionality and follows the user-specified coding style. One model is to provide functionality, and the other is to provide style.
2.1 Task Definition
As Figure 1 shows,given a source code snippet and a reference example that reflects the target style (while maintaining equivalent functionality), the generation process incorporates two modalities: (1) a textual representation of the reference code. (2) a style vector extracted by a dedicated style encoding module.
During training, the model learns to align these modalities such that it can later generate a stylistically consistent version of the source code that mirrors the style characteristics of the reference. This formulation enables fine-grained and continuous control over various dimensions of code style, including formatting, indentation, naming conventions, and structural organization, all without compromising semantic correctness.
2.2 Proposed Method
To overcome these limitations, we propose a novel two-stage framework for style-controllable code generation. Our framework extracts explicit style representations from target-style reference code and uses these style vectors to guide a fine-tuned language model in generating functionally equivalent yet stylistically aligned code.
Given a source code snippet and a reference code sample (with similar functionality but desired style), we first extract a quantitative style vector from the reference code using a style encoder. This style vector captures latent stylistic features such as naming conventions, indentation patterns, whitespace density, and structural layout. The extracted vector is then fused with the source code input and passed into a decoder-style LLM for generation.
As shown in 2, the style vector is first extracted as a 34-dimensional representation, which is then transformed by the style encoder to match the required dimensionality, enabling functional code mixing.
Our model architecture consists of three core components: Style Representation Module This module includes: A style feature extractor that parses source and object code to compute a 34-dimensional interpretable style vector covering aspects such as indentation patterns, loop nesting, whitespace usage, and naming conventions. A style encoder: a multi-layer perceptron (MLP) that transforms the style vector into a dense style embedding space. Language Backbone We use the encoder-decoder language model Flan-T5 as the base generator. During training, its encoder is frozen to ensure semantic stability, while the decoder is fine-tuned to learn style-aware code generation using the injected style embeddings. Style2 extracts code snippets from historical conversations and then passes them on based on whether there are any historical codes before. If they exist, they are superimposed, otherwise they are passed on normally. The passed-in model is mixed with code tokens through the style analyzer and style encoding for style generation.
2.3 Style Analyzer
The Style Analyzer is a key module in our system, designed to explicitly extract structured style vectors from code to support controllable and personalized code generation. It integrates a small number of QA-based vectors with large-scale structural vector analysis and organizes style modeling into three dimensions: Naming Style, Layout Style, and Structural / Functional Style.
2.3.1 Naming Style
Feature | Description |
name_length | Avg name length |
is_snake_case | Is name snake_case |
style_stat_underscore_ratio | Underscore ratio |
style_stat_digit_ratio | Digits in names |
[…]
style_stat_uppercase_ratio | Uppercase char ratio |
style_stat_lowercase_ratio | Lowercase char ratio |
style_dist_PascalCase | PascalCase names |
style_dist_snake_case | snake_case names |
style_dist_camelCase | camelCase names |
style_dist_UPPER_CASE | UPPER_CASE names |
style_dist_private | _private names |
style_dist_dunder_method | __dunder__ methods |
Table 1 describes the layout-related features extracted for style encoding. This part focuses on analyzing variable and function naming conventions. It captures patterns and statistical features that reflect a developer’s naming preference.
2.3.2 Structural and Functional Style
Feature | Description |
call_depth | Max call nesting |
branch_count | Number of branches |
return_count | Number of return statements |
arg_count | Number of arguments |
length | Number of lines |
has_docstring | Has docstring or not |
has_try_except | Contains try/except |
exception_score | Specificity of exception handling |
redundancy_ratio | Code redundancy |
annotation_ratio | Parameters with type hints |
control_structures | Number of control structures |
This category involves abstract syntax tree (AST)-based static analysis, used to measure complexity, structural patterns, and function-level behaviors in code. Table 2 lists the features used for analyzing structure and functionality.
2.3.3 Layout Style
Feature | Description |
comment_ratio | Comment lines / total |
type_hint_ratio | Parameters with type hints |
redundancy_ratio | Code redundancy |
space_pattern_code2 | Space pattern matrix (0/1) |
We propose a visual representation called the Code Space Pattern, which transforms code into fixed-size spatial matrices and learns layout-aware embeddings that reflect indentation, spacing, and formatting habits.
2.4 Style Encoder
To encode the extracted code style vector into a high-dimensional latent space compatible with the backbone language model, we design a 4-layer residual neural network. This encoder takes as input a 34-dimensional hand-crafted style feature vector—encompassing naming conventions, structural patterns, and layout styles—and projects it into a 1024-dimensional latent representation. The primary purpose of this module is to enable the model to recognize and incorporate stylistic modality during contrastive learning, thereby facilitating the alignment between style and code modalities.
3 Training
We employ a publicly available Python code generation dataset containing approximately 18K code samples. To ensure stylistic diversity, we use three state-of-the-art code generation models—DeepSeek, Doubao, and Qwen—to generate stylistic variants of the same semantic function. These variants form pairs with shared semantics but differing styles, each annotated with a 25-dimensional style vector.
3.1 Contrastive Learning
In personalized code generation, understanding code style is a challenging task due to its latent and implicit nature. Unlike functional semantics, which can be directly inferred from token sequences or abstract syntax trees (ASTs), code style emerges from a combination of naming patterns, formatting practices, structural choices, and spacing conventions.Therefore, rather than relying on supervised labels or handcrafted style categories, we use a self-supervised contrastive learning strategy to allow the model to discover style patterns automatically from raw code corpora. The first stage of our framework focuses on learning an explicit and discriminative representation of code style via contrastive learning. This component, named the Style Encoder, is responsible for mapping reference code fragments into a latent style space, where stylistically similar samples are embedded closely, while dissimilar ones are pushed apart.
3.1.1 Positive and Negative Pair Construction
To train the Style Encoder, we construct contrastive pairs from large-scale unlabeled code data. Specifically:Positive pairs are formed by taking a complete code snippet and extracting a style fragment from it (e.g., a function definition, a loop body, or a comment-annotated segment). These fragments naturally share the same underlying style. Negative pairs are formed by pairing a code snippet with style fragments sampled from different files or authors to ensure stylistic dissimilarity.
3.1.2 Loss Function
We adopt the InfoNCE loss function to train the encoder, defined as:
where and are the embeddings of the anchor code and its positive style fragment, are negative samples within the same batch, is the cosine similarity function, and is a temperature hyperparameter.
The Style Encoder is implemented as a lightweight BiGRU network followed by a linear projection layer. During training, it learns to cluster code fragments with similar styles in the embedding space.
3.1.3 Training details
The contrastive learning process is conducted independently of the language model. The language model is frozen in this stage to ensure that the learned style representations are general and not entangled with token-level generation patterns. We train the Style Encoder for 30 epochs using a batch size of 128, temperature , and an embedding dimension of 1024. After training, the encoder is fixed and used in the second stage for style-guided generation.
3.1.4 Joint Fine-tuning with Style Supervision
In the second stage, we train a decoder-only generation model called StyleControlledGenerator, which takes as input a source code snippet (code1) and a target style vector extracted from a reference. The model is supervised using both semantic loss and style loss.
Semantic Loss.
The primary objective is to generate the target code (code2) that matches the ground-truth tokens. We apply the standard token-level cross-entropy loss:
where denotes the input code and style vector, and is the -th token in code2.
Style Loss.
To further ensure that the generated code conforms to the intended style, we use a frozen Style Encoder to extract the style vector from the generated code. We then compute the mean squared error (MSE) between this predicted style vector and the original target style vector:
Joint Objective.
The total loss is a weighted sum of semantic and style objectives:
where is a hyperparameter controlling the influence of style supervision.
4 Experiment
We used 7 sets of model experiments and 4 evaluations to verify the overall performance of our model.
4.1 Evaluation Metrics
To comprehensively assess the quality and stylistic fidelity of generated code, we adopt both standard benchmark metrics and custom-designed indicators. The evaluation framework includes the following:
BLEU-4 (from MPCoder): This metric evaluates the -gram overlap between the generated code and the reference code, focusing on token-level syntactic similarity. A higher BLEU-4 score indicates a closer lexical match.
ROUGE (from MPCoder): We adopt three variants of the ROUGE metric to capture different aspects of similarity:
ROUGE-1: Measures unigram (single-word) overlap, reflecting basic lexical similarity.
ROUGE-2: Measures bigram (two-word sequence) overlap, capturing local phrase-level consistency.
ROUGE-L: Evaluates the longest common subsequence (LCS) between the generated and reference code, serving as a proxy for structural alignment.
CSS (Code Style Similarity): A task-specific metric designed to quantify stylistic alignment between the generated and reference code. It operates on a 14-dimensional vector extracted from surface-level style features, including formatting, naming, and structural conventions.
AI Evaluation: An automatic scoring mechanism powered by a language model-based evaluator. This metric assesses both the semantic correctness and stylistic fidelity of the generated code, simulating a human-like judgment process.
4.2 Quantitative Evaluation Results
To assess the performance of our model Style2Code, we compare it against baseline (Flan-T5) and state-of-the-art models (e.g., DeepSeek, Qwen) using four key evaluation metrics: BLEU-4, AI Judgment, ROUGE, and CSS Similarity. The corresponding results are illustrated in Figures 3, 4, 5, and 6.
BLEU-4 Score.
As illustrated in Figure 3, our model achieves a BLEU-4 score of 0.310, representing a remarkable 1450% improvement over the baseline Flan-T5 (0.020), and an 8.8% gain over the best-performing DeepSeek (code2style1), which achieved a score of 0.285.
AI-Based Style Judgment.
In the evaluation of stylistic consistency by an AI-based style classifier (Figure 4), Style2Code obtains a score of 0.090, significantly surpassing Flan-T5 (close to 0), and ranking second overall. Although trailing behind DeepSeek (0.230), the relative performance gap is approximately 60.9%.
ROUGE F1 Scores.
As shown in Figure 5, Style2Code achieves competitive ROUGE scores across all sub-metrics:
ROUGE-1: 0.330 — a +371% increase over Flan-T5, and on par with the top-performing model.
ROUGE-2: 0.180 — a +350% improvement over Flan-T5, trailing the best model by approximately 18%.
ROUGE-L: 0.240 — a +71% increase compared to Flan-T5, and 7.7% lower than the top result.
CSS (Code Style Similarity).
To assess stylistic alignment, we employ our custom metric CSS, shown in Figure 6. Style2Code achieves the highest CSS score of 0.910, outperforming all baseline models. The strongest baseline, Qwen (code2style1), achieves 0.810, marking a 12.3% relative improvement. These results confirm Style2Code’s superior capacity to preserve and enforce stylistic fidelity in code generation.
5 Conclusion
In this work, we propose a novel style-guided code generation framework that enables the generation of source code conforming to a specified target style. This framework addresses the growing demand for personalized and stylistically consistent code in real-world development environments.
Our key innovation lies in the use of contrastive learning to extract explicit style features from code. The experimental results demonstrate that this approach is effective in capturing fine-grained stylistic characteristics and enhancing the controllability of code generation.
Through extensive evaluation, we show that our model outperforms the baseline in multiple metrics, particularly in text-based and fine-grained stylistic matching. Nonetheless, due to the lack of targeted learning strategies for implicit structural features, our model still falls slightly short of state-of-the-art (SOTA) models in structure-aware metrics.
An important strength of our approach is its complementary effect: while large language models typically excel in generating structurally sound code through implicit prompting, our model supplements this by enhancing fine-grained style control via explicit guidance, forming a more complete solution for code style alignment.
In addition, our 7B-scale implementation demonstrates strong performance, underscoring the scalability and practical effectiveness of the proposed method in large-model settings. By treating code style as an explicit modality, our framework offers a promising direction in the field of multimodal learning, where conditioning on fine-grained, learnable modalities remains a key challenge. This formulation not only improves controllability in code generation, but also introduces a generalizable approach for modality-aware code synthesis.
Finally, we note that the field of code style transfer remains underexplored. Existing methods are mostly confined to the AGENT-level or rely on multimodal strategies for the same content. Our unified, single-model architecture represents a new theoretical direction for this task, contributing to the advancement of both methodology and application in this domain.

### v3

@@
-Style2Code: A Style-Controllable Code Generation Framework with Dual-Modal Contrastive Representation Learning Style2Code: A Style-Controllable Code Generation Framework with Dual-Modal Contrastive Representation Learning
+Style2Code: A Style-Controllable Code Generation Framework with Dual-Modal Contrastive Representation Learning
 Abstract
@@
 To the best of our knowledge, Style2Code is the first model to achieve explicit and user-controllable code style transfer through continuous vector conditioning in a dual-modal generation setting.
-The source code and data sets are publicly available to facilitate further research:
-Style2Code: A Style-Controllable Code Generation Framework with
-Dual-Modal Contrastive Representation Learning
-1 Introduction
-1.1 Problem
+The source code and data sets are publicly available to facilitate further research: https://github.com/zh19980811/Style2Code
+I Introduction
+I-A Problem
 In the daily use of AI models for code generation, it is often observed that the generated code may not fully meet the user’s requirements in terms of style due to the influence of training data and user prompts.
 Additionally, the code style produced by different models may also vary slightly.
-Without providing examples of similar types of code, specifying a particular style in the prompt, or providing detailed algorithm requirements, the model’s output often contains an element of randomness.
-We can categorize the style variation issues in code-generative AI into several types: (1) differences in style output by the same model for identical prompts at different times; (2) style variations between different models at the same time for identical prompts; (3) differences in output from the same model at the same time for prompts with similar structure but different style requirements.
-The code style itself also encompasses multiple aspects: (1) naming differences, such as variations in variable names; (2) functional differences, where the same code may implement functionality in varied ways; (3) structural or functional discrepancies arising from a lack of context, leading to incomplete structure or function implementation.
-The same prompt can lead to different code styles when generated by different models; In simple terms,Even for the same model, repeated generations may show slight stylistic variations (despite similar structure).
-1.2 Background
-Personalized code generation has become an important area in automated code synthesis.
-MPCoder advances this field by incorporating both explicit and implicit style representations for multi-user scenarios.
-However, it faces key limitations: it cannot effectively generalize implicit stylistic cues embedded in code semantics and relies on fixed style encodings and handcrafted metrics, limiting adaptability.
-Additionally, MPCoder demands large user-specific datasets, creating scalability issues—each new user requires retraining, leading to cold-start problems and high maintenance costs.
-This hinders practical deployment in dynamic environments with limited training data.
-While works like Choi et al.
-(2023) explore stylistic clustering for authorship attribution using LLMs, they focus on identification, not controllable generation.
-Thus, there remains a clear need for a scalable, generalizable framework that enables fine-grained, user-controllable code style generation with minimal data and no retraining.
-1.3 Contributions
-Our main contributions are summarized as follows:
+Without providing examples of similar types of code, specifying a particular style in the prompt, or providing detailed algorithm requirements, the model’s output often contains an element of randomness [1].
+In practical software development, inconsistencies between AI-generated code styles and established project conventions can lead to a range of negative consequences.
+First, inconsistent naming conventions force developers to memorize additional variations, thereby increasing cognitive load and reducing team collaboration efficiency.[2] Second, differences in indentation and formatting divert attention during code review away from logical correctness, requiring extra effort for style corrections and often introducing conflicts during pull request integration.[3] Finally, AI-specific code smells, recurring patterns in code that may indicate deeper issues such as irreproducibility, silent failures, or poor model generalization, are often overlooked when users use AI-generated code.
+inconsistencies, which increases both learning and maintenance costs.
+[4]By allowing users to steer AI code generation with specific conditions, the resulting code can better align with project requirements and coding standards, which in turn reduces the effort required for subsequent code maintenance.
+I-B Motivation and Significance
+Our primary goal is to enable AI models to generate code that aligns with the user’s individual or project-specific coding style during the generation process.
+The key component behind this capability is code style—the consistent and recognizable patterns that define how code is written and structured.
+These patterns include indentation schemes, naming conventions, common structural layouts, and preferred utility functions that reflect developers’ habitual practices.
+By adhering to these patterns, AI-generated code becomes more readable, maintainable, and seamlessly integrated into existing codebases.
+In professional development environments, such stylistic uniformity is not merely aesthetic but functional—it improves collaboration, reduces code review costs, and preserves the overall consistency of large software projects.
+Established conventions such as the Google Style Guide, PEP8 for Python, or Airbnb’s JavaScript Style Guide demonstrate that style guidance helps all contributors operate under the same set of assumptions.
+Inspired by these principles, our approach aims to encode and reproduce such stylistic regularities through style-aware representation learning, ensuring that the generated code not only functions correctly but also conforms to the user’s established style.
+This alignment bridges the gap between machine-generated and human-written code, making AI-assisted programming more natural, efficient, and trustworthy.
+II Related Work
+II-A Personalized Code Generation
+Personalized code generation has become an important research direction in automated code synthesis.
+MPCoder [5] represents a notable advance by incorporating both explicit and implicit style representations for multi-user scenarios.
+However, it faces several limitations: (1) difficulty in generalizing implicit stylistic cues embedded within code semantics; (2) reliance on fixed, handcrafted style encodings and metrics, which restrict adaptability; and (3) the requirement of large user-specific datasets, leading to cold-start and scalability issues.
+These constraints hinder its deployment in dynamic environments with limited or evolving data.
+In contrast, our framework aims to achieve style control through transferable representation learning without requiring user-specific retraining.
+II-B Code Style Representation and Authorship Attribution
+Several studies have attempted to model stylistic information in code through language-based or authorship-aware learning.
+Ahmed et al.
+[6] enhance LLMs’ style representation by integrating semantic information via code summarization.
+Similarly, Choi and Tan [7] explore source code authorship attribution using large language models, highlighting its applications in software forensics and plagiarism detection.
+Their findings demonstrate that LLMs can capture consistent stylistic patterns across different programming languages, achieving competitive authorship classification performance under zero- and few-shot prompting.
+Although these works focus on authorship recognition rather than code generation, they suggest that modern LLMs inherently encode stylistic features that can be leveraged for personalized generation.
+II-C Contrastive and Style-Aware Learning
+Recent progress in contrastive learning further motivates our approach.
+Bui et al.
+[8] proposed a semi-supervised contrastive framework that trains source-code encoders to distinguish semantically similar and dissimilar code snippets, enhancing their understanding of structural and functional relationships.
+Building on this, Liu and Nguyen et al.
+[9] employed an iterative synthesis-and-questioning approach to refine code style representations, improving model adaptability and stylistic coherence.
+Unlike these methods, our proposed Style2Code framework performs dual-modality contrastive learning between style and code embeddings, enabling controllable style alignment without requiring extensive labeled data.
+Building upon these insights, the following section introduces the overall architecture and training objectives of Style2Code, which integrate style representation learning with code generation in a unified framework.
+II-D Contributions
+Our work introduces the issue of cross-modal contrastive learning.
+In the analysis of the same problem, we outperform the state-of-the-art model, MPCoder.[5] Through contrastive learning, we enable mutual recognition between code and its style encoding, addressing the cold-start problem.
+Additionally, we tackle the challenge of clustering different code samples.Our main contributions are summarized as follows:
 (1) We design a fine-grained explicit style encoder that directly extracts interpretable stylistic features—such as indentation schemes, whitespace alignment, loop nesting depth, and structural repetition patterns—from raw source code, without relying on manual annotations or handcrafted templates.
-(2) We introduce a contrastive alignment framework that semantically aligns code embeddings with their corresponding style vectors, enabling the model to generalize to unseen styles while preserving controllability during generation.
+(2) We introduce a contrastive alignment framework that semantically aligns code embeddings with their corresponding style vectors, enabling the model to generalize to unseen styles while preserving controllability during generation.Which is very similar to CLIP, but it actually combines learning style and code.[10]
 (3) We propose a two-stage training paradigm that first trains the style encoder independently to obtain stable and accurate style representations, and then injects these representations into the code generator for style-controlled generation.
@@
 (4) Furthermore, we formulate the style transfer task as a multimodal learning problem, and explore its potential in the context of personalized code generation, enabling flexible and user-controllable stylistic variation.
-2 Method
-We define the personalized code generation task as amultimodal generation problem, where the goal is to generate code that both retains the original functionality and follows the user-specified coding style.
-One model is to provide functionality, and the other is to provide style.
-2.1 Task Definition
-As Figure 1 shows,given a source code snippet and a reference example that reflects the target style (while maintaining equivalent functionality), the generation process incorporates two modalities: (1) a textual representation of the reference code.
-(2) a style vector extracted by a dedicated style encoding module.
-During training, the model learns to align these modalities such that it can later generate a stylistically consistent version of the source code that mirrors the style characteristics of the reference.
-This formulation enables fine-grained and continuous control over various dimensions of code style, including formatting, indentation, naming conventions, and structural organization, all without compromising semantic correctness.
-2.2 Proposed Method
-To overcome these limitations, we propose a novel two-stage framework for style-controllable code generation.
-Our framework extracts explicit style representations from target-style reference code and uses these style vectors to guide a fine-tuned language model in generating functionally equivalent yet stylistically aligned code.
-Given a source code snippet and a reference code sample (with similar functionality but desired style), we first extract a quantitative style vector from the reference code using a style encoder.
-This style vector captures latent stylistic features such as naming conventions, indentation patterns, whitespace density, and structural layout.
-The extracted vector is then fused with the source code input and passed into a decoder-style LLM for generation.
-As shown in 2, the style vector is first extracted as a 34-dimensional representation, which is then transformed by the style encoder to match the required dimensionality, enabling functional code mixing.
-Our model architecture consists of three core components: Style Representation Module This module includes: A style feature extractor that parses source and object code to compute a 34-dimensional interpretable style vector covering aspects such as indentation patterns, loop nesting, whitespace usage, and naming conventions.
-A style encoder: a multi-layer perceptron (MLP) that transforms the style vector into a dense style embedding space.
-Language Backbone We use the encoder-decoder language model Flan-T5 as the base generator.
-During training, its encoder is frozen to ensure semantic stability, while the decoder is fine-tuned to learn style-aware code generation using the injected style embeddings.
-Style2 extracts code snippets from historical conversations and then passes them on based on whether there are any historical codes before.
-If they exist, they are superimposed, otherwise they are passed on normally.
-The passed-in model is mixed with code tokens through the style analyzer and style encoding for style generation.
-2.3 Style Analyzer
-The Style Analyzer is a key module in our system, designed to explicitly extract structured style vectors from code to support controllable and personalized code generation.
-It integrates a small number of QA-based vectors with large-scale structural vector analysis and organizes style modeling into three dimensions: Naming Style, Layout Style, and Structural / Functional Style.
-2.3.1 Naming Style
+III Method
+We define the personalized code generation task as a multimodal generation problem, where the goal is to generate code that retains the original functionality while adhering to the user-specified coding style.
+One modality provides functional semantics, and the other provides stylistic guidance [11].
+III-A Task Definition
+As shown in Fig.
+1, given a source code snippet and a reference example that reflects the target style (while maintaining equivalent functionality), the generation process incorporates two modalities: (1) a textual representation of the reference code, and (2) a quantitative style vector extracted by a dedicated style encoding module.
+During training, the model learns to align these modalities such that it can later generate a stylistically consistent version of the source code that mirrors the characteristics of the reference.
+This formulation enables fine-grained and continuous control [12] over various aspects of coding style—such as formatting, indentation, naming conventions, and structural organization—without compromising semantic correctness.
+III-B Proposed Framework
+To overcome the limitations of existing methods, we propose a novel two-stage framework for style-controllable code generation (Fig.
+2).
+The framework explicitly extracts style representations from target-style reference code and uses these representations to guide a fine-tuned language model in generating functionally equivalent yet stylistically aligned code.
+Given a source code snippet and a reference code sample (with similar functionality but desired style), we first extract a 34-dimensional handcrafted style vector from the reference code using a Style Analyzer and encode it through a Style Encoder.
+The resulting embedding captures latent stylistic features such as naming conventions, indentation patterns, whitespace density, and layout organization.
+This encoded vector is fused with the source code input and passed to a decoder-style LLM for generation [13].
+III-C Style Analyzer
+The Style Analyzer is designed to explicitly extract structured and interpretable features that describe a developer’s habitual coding patterns.
+It organizes style modeling into three dimensions: Naming Style, Layout Style, and Structural/Functional Style.
+These extracted features form a 34-dimensional vector used as the input to the style encoder.
+III-C1 Naming Style
 | Feature | Description |
-name_length | Avg name length |
-is_snake_case | Is name snake_case |
-style_stat_underscore_ratio | Underscore ratio |
-style_stat_digit_ratio | Digits in names |
+| name_length | Average name length |
+| is_snake_case | Whether identifier uses snake_case |
+| style_stat_underscore_ratio | Underscore usage ratio |
+| style_stat_digit_ratio | Digits in identifiers |
 […]
-style_stat_uppercase_ratio | Uppercase char ratio |
-style_stat_lowercase_ratio | Lowercase char ratio |
-style_dist_PascalCase | PascalCase names |
-style_dist_snake_case | snake_case names |
-style_dist_camelCase | camelCase names |
-style_dist_UPPER_CASE | UPPER_CASE names |
-style_dist_private | _private names |
-style_dist_dunder_method | __dunder__ methods |
-Table 1 describes the layout-related features extracted for style encoding.
-This part focuses on analyzing variable and function naming conventions.
-It captures patterns and statistical features that reflect a developer’s naming preference.
-2.3.2 Structural and Functional Style
+| style_stat_uppercase_ratio | Uppercase letter ratio |
+| style_stat_lowercase_ratio | Lowercase letter ratio |
+| style_dist_PascalCase | Frequency of PascalCase names |
+| style_dist_snake_case | Frequency of snake_case names |
+| style_dist_camelCase | Frequency of camelCase names |
+| style_dist_UPPER_CASE | Frequency of UPPER_CASE names |
+| style_dist_private | Private variable prefix (_) usage |
+| style_dist_dunder_method | Dunder methods (__init__) usage |
+Table I summarizes the naming-related features used for style encoding.
+These indicators capture both statistical and categorical naming preferences that reflect an individual developer’s or project’s style.
+III-C2 Structural and Functional Style
 | Feature | Description |
-call_depth | Max call nesting |
+| call_depth | Maximum function call depth |
 | branch_count | Number of branches |
 | return_count | Number of return statements |
-arg_count | Number of arguments |
-length | Number of lines |
-has_docstring | Has docstring or not |
-has_try_except | Contains try/except |
+| arg_count | Number of function arguments |
+| length | Lines of code per function |
+| has_docstring | Whether a docstring exists |
+| has_try_except | Presence of try/except blocks |
 | exception_score | Specificity of exception handling |
-redundancy_ratio | Code redundancy |
+| redundancy_ratio | Duplicate code ratio |
 | annotation_ratio | Parameters with type hints |
 | control_structures | Number of control structures |
-This category involves abstract syntax tree (AST)-based static analysis, used to measure complexity, structural patterns, and function-level behaviors in code.
-Table 2 lists the features used for analyzing structure and functionality.
-2.3.3 Layout Style
+This group of features is extracted through abstract syntax tree (AST)-based static analysis, quantifying complexity and structural organization of code [15].
+These indicators capture the developer’s function-level and control-flow preferences.
+III-C3 Layout Style
 | Feature | Description |
-comment_ratio | Comment lines / total |
-type_hint_ratio | Parameters with type hints |
-redundancy_ratio | Code redundancy |
-space_pattern_code2 | Space pattern matrix (0/1) |
-We propose a visual representation called the Code Space Pattern, which transforms code into fixed-size spatial matrices and learns layout-aware embeddings that reflect indentation, spacing, and formatting habits.
-2.4 Style Encoder
-To encode the extracted code style vector into a high-dimensional latent space compatible with the backbone language model, we design a 4-layer residual neural network.
-This encoder takes as input a 34-dimensional hand-crafted style feature vector—encompassing naming conventions, structural patterns, and layout styles—and projects it into a 1024-dimensional latent representation.
-The primary purpose of this module is to enable the model to recognize and incorporate stylistic modality during contrastive learning, thereby facilitating the alignment between style and code modalities.
-3 Training
-We employ a publicly available Python code generation dataset containing approximately 18K code samples.
-To ensure stylistic diversity, we use three state-of-the-art code generation models—DeepSeek, Doubao, and Qwen—to generate stylistic variants of the same semantic function.
+| comment_ratio | Ratio of comment lines to total lines |
+| type_hint_ratio | Parameters with explicit type hints |
+| redundancy_ratio | Code redundancy ratio |
+| space_pattern_code | Encoded space/indentation pattern |
+The layout style focuses on whitespace, indentation, and comment usage.
+We define a visual descriptor called the Code Space Pattern, which converts code into a fixed-size binary matrix representing indentation and spacing regularities.
+This captures the developer’s layout habits that contribute to visual readability.
+III-D Style Encoder
+To encode the extracted style vector into a high-dimensional latent space compatible with the backbone language model, we employ a four-layer multilayer perceptron (MLP) with one residual connection between the second and fourth layers.
+This design balances representation capacity and training stability.
+Formally, given a 34-dimensional handcrafted feature vector , the style encoder projects it into a 1024-dimensional latent embedding :
+where denotes the residual mapping applied across intermediate layers.
+This embedding serves as the conditioning signal injected into the code generator, allowing the model to align output style with the provided reference.
+During inference, user-provided style reference code samples—referred to as historical dialogues—are analyzed by the Style Analyzer to produce , which is then encoded and injected into the LLM.
+If no prior style reference exists, the model defaults to neutral style generation.
+This mechanism ensures controllable, user-specific stylistic adaptation during the code generation process [16, 17].
+IV Training
+We employ a publicly available Python code generation dataset containing approximately 18K code samples.[19] To ensure stylistic diversity, we use three state-of-the-art code generation models—DeepSeek, Doubao,[18] and Qwen[20]—to generate stylistic variants of the same semantic function.
 These variants form pairs with shared semantics but differing styles, each annotated with a 25-dimensional style vector.
-3.1 Contrastive Learning
+IV-A Contrastive Learning
 In personalized code generation, understanding code style is a challenging task due to its latent and implicit nature.
-Unlike functional semantics, which can be directly inferred from token sequences or abstract syntax trees (ASTs), code style emerges from a combination of naming patterns, formatting practices, structural choices, and spacing conventions.Therefore, rather than relying on supervised labels or handcrafted style categories, we use a self-supervised contrastive learning strategy to allow the model to discover style patterns automatically from raw code corpora.
+Unlike functional semantics, which can be directly inferred from token sequences or abstract syntax trees , code style emerges from a combination of naming patterns, formatting practices, structural choices, and spacing conventions.Therefore, rather than relying on supervised labels or handcrafted style categories, we use a self-supervised contrastive learning strategy to allow the model to discover style patterns automatically from raw code corpora.
 The first stage of our framework focuses on learning an explicit and discriminative representation of code style via contrastive learning.
-This component, named the Style Encoder, is responsible for mapping reference code fragments into a latent style space, where stylistically similar samples are embedded closely, while dissimilar ones are pushed apart.
-3.1.1 Positive and Negative Pair Construction
+This component, named the Style Encoder, is responsible for mapping reference code fragments into a latent style space, where stylistically similar samples are embedded closely, while dissimilar ones are pushed apart.[21]
+IV-A1 Positive and Negative Pair Construction
 To train the Style Encoder, we construct contrastive pairs from large-scale unlabeled code data.
 Specifically:Positive pairs are formed by taking a complete code snippet and extracting a style fragment from it (e.g., a function definition, a loop body, or a comment-annotated segment).
-These fragments naturally share the same underlying style.
-Negative pairs are formed by pairing a code snippet with style fragments sampled from different files or authors to ensure stylistic dissimilarity.
-3.1.2 Loss Function
+These fragments naturally share the same underlying style.[22] Negative pairs are formed by pairing a code snippet with style fragments sampled from different files or authors to ensure stylistic dissimilarity.
+IV-A2 Loss Function
 We adopt the InfoNCE loss function to train the encoder, defined as:
@@
 The Style Encoder is implemented as a lightweight BiGRU network followed by a linear projection layer. During training, it learns to cluster code fragments with similar styles in the embedding space.
-3.1.3 Training details
+IV-A3 Training details
 The contrastive learning process is conducted independently of the language model.
@@
 After training, the encoder is fixed and used in the second stage for style-guided generation.
-3.1.4 Joint Fine-tuning with Style Supervision
+IV-A4 Joint Fine-tuning with Style Supervision
 In the second stage, we train a decoder-only generation model called StyleControlledGenerator, which takes as input a source code snippet (code1) and a target style vector extracted from a reference.
@@
 where is a hyperparameter controlling the influence of style supervision.
-4 Experiment
+V Experiment
 We used 7 sets of model experiments and 4 evaluations to verify the overall performance of our model.
-4.1 Evaluation Metrics
+V-A Evaluation Metrics
 To comprehensively assess the quality and stylistic fidelity of generated code, we adopt both standard benchmark metrics and custom-designed indicators.
@@
 BLEU-4 (from MPCoder): This metric evaluates the -gram overlap between the generated code and the reference code, focusing on token-level syntactic similarity.
-A higher BLEU-4 score indicates a closer lexical match.
-ROUGE (from MPCoder): We adopt three variants of the ROUGE metric to capture different aspects of similarity:
+A higher BLEU-4 score indicates a closer lexical match.[23]
+ROUGE (from MPCoder): We adopt three variants of the ROUGE metric to capture different aspects of similarity[24]:
 ROUGE-1: Measures unigram (single-word) overlap, reflecting basic lexical similarity.
@@
 This metric assesses both the semantic correctness and stylistic fidelity of the generated code, simulating a human-like judgment process.
-4.2 Quantitative Evaluation Results
+V-B Quantitative Evaluation Results
 To assess the performance of our model Style2Code, we compare it against baseline (Flan-T5) and state-of-the-art models (e.g., DeepSeek, Qwen) using four key evaluation metrics: BLEU-4, AI Judgment, ROUGE, and CSS Similarity.
-The corresponding results are illustrated in Figures 3, 4, 5, and 6.
+The corresponding results are illustrated in Fig.
+3, 4, 5, and 6.
 BLEU-4 Score.
-As illustrated in Figure 3, our model achieves a BLEU-4 score of 0.310, representing a remarkable 1450% improvement over the baseline Flan-T5 (0.020), and an 8.8% gain over the best-performing DeepSeek (code2style1), which achieved a score of 0.285.
+As illustrated in Fig.
+3, our model achieves a BLEU-4 score of 0.317, representing a remarkable 1450% improvement over the baseline Flan-T5 (0.020), and an 8.8% gain over the best-performing DeepSeek (code2style1), which achieved a score of 0.285.It is also higher than the SOTA model mpcoder.
 AI-Based Style Judgment.
-In the evaluation of stylistic consistency by an AI-based style classifier (Figure 4), Style2Code obtains a score of 0.090, significantly surpassing Flan-T5 (close to 0), and ranking second overall.
-Although trailing behind DeepSeek (0.230), the relative performance gap is approximately 60.9%.
+In the evaluation of stylistic consistency by an AI-based style classifier (Fig 4), Style2Code obtains a score of 0.115, significantly surpassing Flan-T5 (close to 0), and ranking third overall.
 ROUGE F1 Scores.
-As shown in Figure 5, Style2Code achieves competitive ROUGE scores across all sub-metrics:
-ROUGE-1: 0.330 — a +371% increase over Flan-T5, and on par with the top-performing model.
-ROUGE-2: 0.180 — a +350% improvement over Flan-T5, trailing the best model by approximately 18%.
-ROUGE-L: 0.240 — a +71% increase compared to Flan-T5, and 7.7% lower than the top result.
+Style2Code achieves competitive ROUGE scores across all sub-metrics: As shown in Fig. 5,Compared to the second-best model, MPCoder, Style2Code achieves higher scores across all three ROUGE metrics:
+ROUGE-1: Style2Code achieves 0.448, outperforming MPCoder (0.371) by +20.8%.
+ROUGE-2: Style2Code scores 0.167, representing a +16.8% improvement over MPCoder (0.143).
+ROUGE-L: Style2Code reaches 0.401, slightly higher than MPCoder (0.388), with a margin of +3.4%.
+Style2Code ranks 1st among all evaluated models on all ROUGE sub-metrics, demonstrating its strong advantage in stylistic alignment and content similarity.
 CSS (Code Style Similarity).
-To assess stylistic alignment, we employ our custom metric CSS, shown in Figure 6.
-Style2Code achieves the highest CSS score of 0.910, outperforming all baseline models.
-The strongest baseline, Qwen (code2style1), achieves 0.810, marking a 12.3% relative improvement.
-These results confirm Style2Code’s superior capacity to preserve and enforce stylistic fidelity in code generation.
-5 Conclusion
+To assess stylistic alignment, we employ our custom metric CSS, as shown in Fig.
+6.
+Style2Code achieves a CSS score of 0.196, ranking third among all evaluated models.
+The top-performing model, Qwen (code2style1), reaches 0.222, followed closely by MPCoder (0.214), which serves as a strong SOTA reference in style-conditioned code generation.
+Compared to MPCoder, Style2Code shows a relative gap of 8.4%, yet achieves comparable performance without explicit multi-task pretraining.
+This highlights Style2Code’s competitive stylistic control in a more lightweight setting.
+In contrast, Style2Code outperforms the baseline model Flan-T5 (0.100) by a substantial margin of +96%, clearly demonstrating the advantage of incorporating style-specific conditioning.
+The results confirm Style2Code’s effectiveness in enforcing stylistic fidelity while maintaining generalization across examples.
+VI Conclusion
 In this work, we propose a novel style-guided code generation framework that enables the generation of source code conforming to a specified target style.